OpenStudy (anonymous):

Combine Fraction (2m/t)+(5/mt)

OpenStudy (anonymous):

Get a common denominator: \[\frac{2m*m}{m*t}+\frac{5}{mt}=\frac{2m^2+5}{mt}\]

OpenStudy (anonymous):

thanks